OpenStudy (anonymous):

factor the sum of cubes y^3+64

OpenStudy (akshay_budhkar):

\[a^3+b^3= (a+b)(a^2-ab+b^2)\]

OpenStudy (anonymous):

y^3 + 4^3 = y^3 + 64 (y+4) (y^2 - 4y + 16)

OpenStudy (akshay_budhkar):

so here we have \[a=y,b=4\]

OpenStudy (anonymous):

(y+4) (y^2-4 y+16)